# Classification of superpotentials for cohomogeneity one Ricci solitons

### v1

@@ 5.1. Singular initial value problem
By direct computation, the first-order system (5.1) takes the form
(5.2) $\begin{split}\dot{q}_{1}=&\,\sqrt{-2A}e^{\frac{q_{1}}{2}-q_{2}}-\sqrt{-2A}e^{\frac{q_{1}}{2}-q_{3}}\\ \dot{q}_{2}=&\,-\frac{\sqrt{E}}{2}e^{\frac{q_{2}}{2}-\frac{q_{3}}{2}}+\frac{\sqrt{E}}{2}e^{-\frac{q_{2}}{2}+\frac{q_{3}}{2}}+\frac{\lambda}{\sqrt{E}}e^{-\frac{q_{2}}{2}-\frac{q_{3}}{2}}-\sqrt{-2A}e^{\frac{q_{1}}{2}-q_{2}}\\ \dot{q}_{3}=&\,\frac{\sqrt{E}}{2}e^{\frac{q_{2}}{2}-\frac{q_{3}}{2}}-\frac{\sqrt{E}}{2}e^{-\frac{q_{2}}{2}+\frac{q_{3}}{2}}+\frac{\lambda}{\sqrt{E}}e^{-\frac{q_{2}}{2}-\frac{q_{3}}{2}}+\sqrt{-2A}e^{\frac{q_{1}}{2}-q_{3}},\\ \dot{u}=&\,-\frac{\sqrt{E}}{2}e^{\frac{q_{2}}{2}-\frac{q_{3}}{2}}-\frac{\sqrt{E}}{2}e^{-\frac{q_{2}}{2}+\frac{q_{3}}{2}}+\frac{\lambda}{\sqrt{E}}e^{-\frac{q_{2}}{2}-\frac{q_{3}}{2}}.\end{split}$
@@
Similarly to [7], we define the new radial coordinate $ds=\sqrt{\alpha}\,dt$, and sometimes denote $\frac{d}{ds}$ by . Then (5.3) becomes
@@
(5.4) $\begin{split}\beta_{1}^{\prime}&=\frac{1}{s}\left(\frac{E}{2\lambda}\beta_{1}(\beta_{2}-\beta_{1})+\beta_{1}\right)-1\\ \beta_{2}^{\prime}&=\frac{1}{s}\left(-\frac{E}{2\lambda}\beta_{2}(\beta_{2}-\beta_{1})+\beta_{2}\right)+1.\end{split}$
This is a singular initial value problem for which we will apply the criteria in [9] to find solutions. Let $\beta_{1}^{0}=\beta_{1}(s=0)$ and $\beta_{2}^{0}=\beta_{2}(s=0)$. Then the condition (i) of [9, Theorem 4.7] is
@@ 5.2. Boundary conditions and complete soliton
$f(t)=\frac{4}{\sqrt{-2AE}}\left(1+\frac{8}{Es(t)}\right)^{-\frac{1}{2}},\qquad g_{1}(t)=\sqrt{s(t)},\qquad g_{2}(t)=\sqrt{s(t)+\frac{8}{E}},$
@@
To summarise, let $V_{i}$, $i=1,2$ be two copies of $S^{2}$, with metrics $h_{i}=\frac{1}{\sqrt{2}}g_{S^{2}}$, where $g_{S^{2}}$ is the round metric. Let $\pi_{i}:V_{1}\times V_{2}\rightarrow V_{i}$ be the projection onto the $i$-th factor. Let $a_{i}\in H^{2}(V_{i},\mathbb{Z})$ be indivisible cohomology classes. Let $P_{q}$ denote the principal $U(1)$-bundle over $V_{1}\times V_{2}$ with Euler class $b_{1}\pi_{1}^{*}a_{1}+b_{2}\pi_{2}^{*}a_{2}$. Since the boundary conditions require $A=-1/2=-\frac{1}{4}d_{i}b_{i}^{2}=-b_{i}^{2}/2$ [2], we may take $b_{1}=b_{2}=-1$ as in [7], so that the Euler class of $P_{q}$ is $-\pi_{1}^{*}a_{1}-\pi_{2}^{*}a_{2}$. Let $\theta$ be the principal $U(1)$-connection on $P_{q}$ with curvature $\Omega=-\pi_{1}^{*}\eta_{1}-\pi_{2}^{*}\eta_{2}$, where $\eta_{i}$ is the Kähler form of $h_{i}$. Let $g_{t}$ denote the metric on $(0,\infty)\times P_{q}$ given by
$g_{t}=f(t)^{2}\theta\otimes\theta+g_{1}(t)^{2}\pi_{1}^{*}h_{1}+g_{2}(t)^{2}\pi_{2}^{*}h_{2},$
where, for any $E>0$,
@@
With the soliton potential $u(t)=-Es(t)/4$, the manifold $(0,\infty)\times P_{q}$ extends to a smooth steady complete gradient Ricci soliton for each $E>0$. Note that the Euler class is the same as required by [7, Theorem 4.20(i)].
@@ 6. Generalised first integrals
A generalised first integral for the soliton equations is defined as a function $F:M\rightarrow\mathbb{R}$ satisfying
@@
where for each $\mathbf{b}$, $F_{\mathbf{b}}$ and $\Phi_{\mathbf{b}}$ are polynomials in $\mathbf{p}$. Letting $\psi=\Phi-\frac{1}{2}\mathbf{d}\cdot\nabla_{\mathbf{p}}F$, we obtain that (6.1) is equivalent to the recurrence relation shown in [2, (4.2)].
@@ 6.1. Generalised first integrals for the Bryant soliton
Acknowledgements. I am grateful to Andrew Dancer and Jason Lotay for valuable discussions and guidance, and to Andrew Dancer in particular for a careful proofreading of the manuscript.

### v2

@@ 5.1. Singular initial value problem
By direct computation, we see that the first-order system (5.1) takes the form
(5.2) $\begin{split}\dot{q}_{1}=&\,\sqrt{-2A}e^{\frac{q_{1}}{2}-q_{2}}-\sqrt{-2A}e^{\frac{q_{1}}{2}-q_{3}}\\ \dot{q}_{2}=&\,-\frac{\sqrt{E}}{2}e^{\frac{q_{2}}{2}-\frac{q_{3}}{2}}+\frac{\sqrt{E}}{2}e^{-\frac{q_{2}}{2}+\frac{q_{3}}{2}}+\frac{\lambda}{\sqrt{E}}e^{-\frac{q_{2}}{2}-\frac{q_{3}}{2}}-\sqrt{-2A}e^{\frac{q_{1}}{2}-q_{2}}\\ \dot{q}_{3}=&\,\frac{\sqrt{E}}{2}e^{\frac{q_{2}}{2}-\frac{q_{3}}{2}}-\frac{\sqrt{E}}{2}e^{-\frac{q_{2}}{2}+\frac{q_{3}}{2}}+\frac{\lambda}{\sqrt{E}}e^{-\frac{q_{2}}{2}-\frac{q_{3}}{2}}+\sqrt{-2A}e^{\frac{q_{1}}{2}-q_{3}}\\ \dot{u}=&\,-\frac{\sqrt{E}}{2}e^{\frac{q_{2}}{2}-\frac{q_{3}}{2}}-\frac{\sqrt{E}}{2}e^{-\frac{q_{2}}{2}+\frac{q_{3}}{2}}+\frac{\lambda}{\sqrt{E}}e^{-\frac{q_{2}}{2}-\frac{q_{3}}{2}}.\end{split}$
@@
Similarly to [7], we define the new radial coordinate $ds=\sqrt{\alpha}\,dt$. Then (5.3) becomes
@@
(5.4) $\begin{split}\beta_{1}^{\prime}&=\frac{1}{s}\left(\frac{E}{2\lambda}\beta_{1}(\beta_{2}-\beta_{1})+\beta_{1}\right)-1\\ \beta_{2}^{\prime}&=\frac{1}{s}\left(-\frac{E}{2\lambda}\beta_{2}(\beta_{2}-\beta_{1})+\beta_{2}\right)+1,\end{split}$
where denotes $\frac{d}{ds}$. This is a singular initial value problem for which we will apply the criteria in [9] to find solutions. Let $\beta_{1}^{0}=\beta_{1}(s=0)$ and $\beta_{2}^{0}=\beta_{2}(s=0)$. Then condition (i) of [9, Theorem 4.7] is
@@ 5.2. Boundary conditions and complete soliton
$f(t)=\frac{4}{\sqrt{-2AE}}\left(1+\frac{8}{Es(t)}\right)^{-\frac{1}{2}},\qquad g_{1}(t)=\sqrt{s(t)},\qquad g_{2}(t)=\sqrt{s(t)+\frac{8}{E}}$
@@
To summarise, let $V_{i}$, $i=1,2$ be two copies of $S^{2}$, with metrics $h_{i}=\frac{1}{\sqrt{2}}g_{S^{2}}$, where $g_{S^{2}}$ is the standard metric on the unit sphere, and let $\pi_{i}:V_{1}\times V_{2}\rightarrow V_{i}$ be the projection onto the $i$-th factor. Let $a_{i}\in H^{2}(V_{i},\mathbb{Z})$ be indivisible cohomology classes. Finally, let $P_{q}$ denote the principal $U(1)$-bundle over $V_{1}\times V_{2}$ with Euler class $b_{1}\pi_{1}^{*}a_{1}+b_{2}\pi_{2}^{*}a_{2}$. Since the boundary conditions require $A=-1/2=-d_{i}b_{i}^{2}/4=-b_{i}^{2}/2$ [2], we may take $b_{1}=b_{2}=-1$, so that the Euler class of $P_{q}$ is $-\pi_{1}^{*}a_{1}-\pi_{2}^{*}a_{2}$ (which is the same as required by [7, Theorem 4.20(i)]). Let $\theta$ be the principal $U(1)$-connection on $P_{q}$ with curvature $\Omega=-\pi_{1}^{*}\eta_{1}-\pi_{2}^{*}\eta_{2}$, where $\eta_{i}$ is the Kähler form of $h_{i}$.
Proposition 5.1.
For each $E>0$, consider the metric $\hat{g}=dt^{2}+g_{t}$ on $(0,\infty)\times P_{q}$, where
$g_{t}=f(t)^{2}\theta\otimes\theta+g_{1}(t)^{2}\pi_{1}^{*}h_{1}+g_{2}(t)^{2}\pi_{2}^{*}h_{2}$
is the metric on $P_{q}$ given by
@@
With the soliton potential $u(t)=-Es(t)/4$, $\hat{g}$ extends to a smooth steady complete gradient Ricci soliton on the compactification of $(0,\infty)\times P_{q}$ by adding $S^{2}$ at $t=0$.
@@ 6. Generalised first integrals
A generalised first integral for the Ricci soliton equations is defined as a function $F:M\rightarrow\mathbb{R}$ satisfying
@@
where, for each $\mathbf{b}$, $F_{\mathbf{b}}$ and $\Phi_{\mathbf{b}}$ are polynomials in $\mathbf{p}$. Letting $\psi=\Phi-\frac{1}{2}\mathbf{d}\cdot\nabla_{\mathbf{p}}F$, we obtain that (6.1) is equivalent to the recurrence relation shown in [2, (4.2)].
@@ 6.1. Generalised first integrals for the Bryant soliton
Acknowledgements. I am grateful to Andrew Dancer and Jason Lotay for valuable discussions, guidance and comments on the manuscript.